OpenStudy (anonymous):

Hi guys! i have a question about slope intercept form. So I have these coordinates (-8,4) and (-8,2) If I did the math right, the slope should be 0. how do u right this all in slope intercept form. I'm not really sure where to start so if you help me I will become fan and give out a medal! thanks in advance!

OpenStudy (anonymous):

can you help me @Owlcoffee

OpenStudy (owlcoffee):

Yes, i was thinking how I could explain it.

OpenStudy (anonymous):

alright ill wait(:

OpenStudy (anonymous):

also if you could put it into standard form it would really help!!

OpenStudy (anonymous):

y = whatever b is.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

could you explain it?

OpenStudy (anonymous):



OpenStudy (anonymous):

Also can you tell me the standard form and the x and y intercepts

OpenStudy (owlcoffee):

So we have the points, let's call them A and B, where A(-8,4) and B(-8,2). We'll begin by looking at the slope formula, applied to the points in question: \[m=\frac{ y _{b}-y _{a} }{ x _{b}-x _{a} }\] So, you say that replacing the points we get a slope of 0, and let's take a look at the formula that allows us to determine the line equation having a pointand the slope: \[(y-y _{1})=m(x-x _{1})\] I'll take point A, and replace it: \[(y-4)=0(x+8)\] \[(y-4)=0\] \[y=4\] Now, let's look at the point-slope formula: \[y=mx+b\] but you said that the slope is 0, and that's true, so we have a formula in the form of: \[y=b\] That means that the line, is parallel to the x-axis, like this: 

OpenStudy (anonymous):

the standard form is x + y= 4

OpenStudy (anonymous):

What about the intercepts?

OpenStudy (anonymous):

working on it

OpenStudy (anonymous):

0 for the x intercept and 4 for the y

OpenStudy (owlcoffee):

if the slope is 0, then there is no x intercept. The formula isn't y=x+b because then you'll presuming that the slope is 1. And that's not true.

OpenStudy (anonymous):

oops! i meant y = 4

OpenStudy (owlcoffee):

The standard or general form of it is y-4=0

OpenStudy (anonymous):

sorry what? just to clarify what are the intercepts, the slope intercept form and the standard form?

OpenStudy (texaschic101):

slope = (2 - 4) / (-8 - (-8) slope = -2/0 If the 0 is in the denominator, then it is undefined slope, not 0 slope slope intercept form is : x = -8

OpenStudy (anonymous):

slope intercept form is y=4, standard is y+4=0, there is no x intercept, and y intercept is 4

OpenStudy (owlcoffee):

Oh dear...

OpenStudy (anonymous):

THANK YOU EVERYONE!!!

OpenStudy (texaschic101):

answer is x = -8